It is claimed that the mean annual salary of call center customer service representatives






is P188,584.00. A researcher randomly selected 50 call center customer service representatives. He computed the mean of their annual salaries and obtained a mean of P188,600.00.

1
Expert's answer
2022-05-04T12:22:45-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le188584"

"H_a:\mu>188584"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05,"

and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z > 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{188600-188584}{39.5/\sqrt{50}}\approx2.86423"

Since it is observed that "z = 2.86423 >1.6449= z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>2.86423)=0.002090," and since "p = 0.002090<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than "188584," at the "\alpha = 0.05" significance level.